Find the solution of the system: y=x to the second power +5
y=2x+4

Answer by checkley71(8403)

y=x^2+5 now substitute (x^2+5) for y in the other equation.
x^2+5=2x+4
x^2-2x+5-4=0
x^2-2x+1=0
(x-1)(x-1)=0
x-1=0
x=1 answer.
y=1^2+5
y=1+5
y=6 answer.
proof
6=2*1+4
6=2+4
6=6